OpenStudy (anonymous):

Simplify 3(4)^2 Answer (a) 24 (b) 36 (c) 48 (d) 144

OpenStudy (anonymous):

C = 48

OpenStudy (anonymous):

You always start with exponents. So (4)^2 = 16*3= 48. :)

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

Your welcome (: